# Efficient Retrieval with Learned Similarities

### v1

@@ 4.2 Top $K$ retrieval performance
5 Related work
Retrieval in Recommendation Models.
Most modern recommendation models follow a classical two-stage information retrieval paradigm (Covington et al., 2016), where up to billions of items (Eksombatchai et al., 2018; Li et al., 2019) are first filtered down to hundreds to thousands in the retrieval or candidate generation stage, followed by the ranking stage that produces the final top 1-10 candidates. Earlier work on scaling large-scale recommendation models primarily utilize dual-encoder setups(Covington et al., 2016) or sometimes graph algorithms (Eksombatchai et al., 2018). Researchers quickly realized that dot products limited retrieval stage’s performance, and introduced multi-query embeddings (Li et al., 2019) and specialized neural networks (He et al., 2017) to alleviate the expressiveness constraints in retrieval. Learned index structures, or generative retrieval, map each candidate item to a series of tokens and leverage beam search to replace approximate nearest neighbor search. These tokens are represented either implicitly during tree traversal (Zhu et al., 2018; Zhuo et al., 2020) or explicitly (Gao et al., 2021). However, they introduce stability challenges as recommendation systems need to support billion-scale dynamic realtime updated set of items. A latest trend is to exploit characteristics of GPUs, TPUs, and other accelerators; learned similarity functions as well as hardware-aware algorithms have been shown to achieve significant quality improvements (Johnson et al., 2021; Chern et al., 2022; Zhai et al., 2023).
Efficient Nearest Neighbor Search (NNS).
Nearest neighbor search has been a popular topic of research due to their critical role in large-scale retrieval and vector databases. Most studies focus on the dot product case, also known as Maximum Inner Product Search (MIPS). Various techniques were proposed and analyzed, including tree structures (Bentley, 1975; Ram and Gray, 2012), locality sensitive hashing (Gionis et al., 1999; Shrivastava and Li, 2014), production quantization (Jegou et al., 2011), data partitioning (Li et al., 2002; Zhai et al., 2011), graph-based methods (Malkov and Yashunin, 2020; Jayaram Subramanya et al., 2019), and so on. The general case for NNS utilizing learned similarities remains less studied; for learned index structures, techniques to construct trees while maintaining max-heap property have been proposed to ensure beam search result in globally optimal top-$K$ results (Zhuo et al., 2020). Due to GPUs and other accelerators having orders of magnitude higher arithmetic intensity vs CPUs, traditional quantization techniques (Shrivastava and Li, 2014; Guo et al., 2016) no longer fully utilize GPUs; to address this issue, accelerator-specific nearest neighbor algorithms that benefit from increased compute have been proposed and analyzed in recent work (Chern et al., 2022; Zhai et al., 2023; Ootomo et al., 2024).
6 Conclusion
We have analyzed techniques for efficient retrieval with learned similarities in this work. We begin by showing Mixture-of-Logits (MoL) is a universal approximator of learned similarity functions, and empirically MoL consistently outperforms dot products in HR across 6 scenarios by an average of 21.4% in HR@1 and 13.7% in HR@10, setting new state-of-the-art on recommendation datasets. We next propose both exact and approximate algorithms to enable efficient retrieval using learned similarity functions, and show their correctness and bounds. Across all datasets evaluated, we demonstrate that our approximate top $K$ algorithms can reach .99.99.99.99 of Hit Rate relative to exact algorithms, while achieving up to 91×91$\times$91 × reduction in end-to-end latency and with minimal indexing overheads. We expect the speedups to be further amplified with larger-scale datasets and GPU kernel optimizations. We believe this result is of interest to both recommendation system research communities and researchers working on enabling advanced learned similarities in vector databases.

### v2

@@ 6 Conclusion
Appendix A Experiment Setups
Reproducibility.
Our code is publicly available online, and detailed implementations and hyperparameter settings for reproducing our experiment results can be found at the following url: https://github.com/bailuding/rails. We discuss key settings below.
Parameterization of low-rank (“component-level”) embeddings, $f_{p}(q)$ and $g_{p}(x)$s for $1\leq p\leq P$.
We note that embedding parameterization was not explicitly discussed in the original Mixture-of-Logits paper [42]. A specific implementation utilizing cluster id information has been proposed and discussed in a follow up work [3], where in the context of large-scale recommendation systems, the authors suggested cluster information based on interests of cohorts of members and topics of posts by themselves can lead to 10% recall gain at $K=400$. However, we cannot easily access similar information in the publicly available MovieLens and Amazon Reviews datasets. We therefore follow implementation provided by the authors [42] and additionally optionally utilizes a User ID keyed one-hot embedding as one query-side low-rank (“component-level”) embeddings $f_{p}(q)$, which is a widely used technique in recommendation systems [24]. All other $f_{p}(q)$s and $g_{p}(x)$s are obtained by applying an MLP on top of query-side/item-side representations in standard sequential recommendation setups [15, 20].
Parameterization of $\pi_{p}(q,x)$ matrices.
We follow the implementation provided by the authors [42], which parameterizes $\pi_{p}(q,x)$ as a two-layer multi-layer perceptron (MLP) with SiLU [7] non-linearity. The inputs to this MLP consist of user-side features, item-side features, and the $P$ dot products $\langle f_{p}(q),g_{p}(x)\rangle$s between the learned low-rank embeddings.
Hyperparameter settings.
We use an identical number of sampled negatives for dot product baselines (cosine similarity, “SASRec”, “HSTU” rows in Table 3) and Mixture-of-logits (“SASRec + MoL”, “HSTU + MoL” rows in Table 3) to ensure a fair comparison, which is $128$ for ML-1M and ML-20M and $512$ for Amazon Books following prior work. For “+ MoL” rows, we additionally grid searched $|P_{x}|$ in $\{2,4,8,16\}$, $d_{P}$ in $\{32,64,128\}$), whether to enable user-id based learned embeddings, and the dropout rate to apply to user-id based embeddings in $\{0.2,0.5,0.8\}$ for the smaller MovieLens datasets. Due to resource constraints, we followed initial hyperparameters provided by the authors [42] for all other parameters.
A.1 Efficiency
Method HR@10 HR@50 MRR Time/Epoch
ML-1M SASRec .2818 .5470 .1352 37s
SASRec + MoL .3034 (+7.7%) .5655 (+3.4%) .1448 (+7.1%) 44s (+18.9%)
HSTU .3332 .5956 .1579 41s
HSTU + MoL .3460 (+3.8%) .6048 (+1.5%) .1673 (+6.0%) 46s (+12.2%)
ML-20M SASRec .2883 .5484 .1375 197s
SASRec + MoL .3124 (+8.4%) .5685 (+3.7%) .1551 (+12.8%) 449s (+127.9%)
HSTU .3557 .6146 .1800 526s
HSTU + MoL .3697 (+3.9%) .6244 (+1.6%) .1880 (+4.4%) 773s (+47.0%)
Books SASRec .0306 .0754 .0153 506s
SASRec + MoL .0402 (+31.4%) .0970 (+28.6%) .0203 (+32.7%) 964s (+90.5%)
HSTU .0469 .1066 .0233 1883s
HSTU + MoL .0613 (+30.7%) .1292 (+21.2%) .0315 (+35.2%) 2236s (+18.7%)
Table 5: Evaluation of recommendation retrieval performance, utilizing the next-item prediction task. Note that these results are identical to those reported in Table 3 but with relative gains (MoL vs dot product) and training cost changes illustrated under different sequential encoder backbones.
Training efficiency.
We measure the training efficiency of the setups in Section 4 over one epoch of data, utilizing a single RTX 6000 Ada GPU and with a constant batch size of 128 for ML-1M and ML-20M, and 64 for Amazon Books in Table 5. Consistent with prior findings [42, 3], MoL’s model quality increases do not generally come with a large increase in training time, thanks to MoL’s higher arithmetic intensity on GPUs, with an average increase of 53% across the six scenarios. To contextualize this 1.53x training cost, prior work on sequential recommenders, such as self-supervised bidirectional models, came with a 5.9x training time increase on ML-20M and 4.5x training time increase on Amazon Reviews (BERT4Rec vs SASRec rows) [23]. We expect additional GPU kernel-level optimizations, such as kernel fusions, to further improve efficiency of the “+ MoL” rows.
Inference efficiency.
Given the number of items (3K, 24K, and 674K) in these datasets, as discussed in Table 2, these public benchmark datasets are not good candidates for measuring inference latency. We note that prior work have reported batched inference latency between 5ms and 20ms on datasets with hundreds of millions to billions of items [42, 3], which is inline with optimized MIPS baselines.